# OpenFake: An Open Dataset and Platform Toward Large-Scale Deepfake Detection

### v1

@@ Abstract
Deepfakes, synthetic media created using advanced AI techniques, have intensified the spread of misinformation, particularly in politically sensitive contexts. Existing deepfake detection datasets are often limited, relying on outdated generation methods, low realism, or single-face imagery, restricting the effectiveness for general synthetic image detection. By analyzing social media posts, we identify multiple modalities through which deepfakes propagate misinformation. Furthermore, our human perception study demonstrates that recently developed proprietary models produce synthetic images increasingly indistinguishable from real ones, complicating accurate identification by the general public. Consequently, we present a comprehensive, politically-focused dataset specifically crafted for benchmarking detection against modern generative models. This dataset contains three million real images paired with descriptive captions, which are used for generating 963k corresponding high-quality synthetic images from a mix of proprietary and open-source models. Recognizing the continual evolution of generative techniques, we introduce an innovative crowdsourced adversarial platform, where participants are incentivized to generate and submit challenging synthetic images. This ongoing community-driven initiative ensures that deepfake detection methods remain robust and adaptive, proactively safeguarding public discourse from sophisticated misinformation threats.
@@ 1 Introduction
Deepfakes, realistic synthetic media generated by AI, have emerged as a serious threat to the information ecosystem (Canadian Security Intelligence Service, 2023; Bengio et al., 2025c). By enabling anyone to fabricate audio-visual content of real people, deepfakes can spread false information at an unprecedented scale, eroding trust across various platforms, from social media and online content to traditional media outlets. High-profile cases (e.g., forged speeches or imagery of public figures) and the prevalence of non-consensual intimate imagery underscore the potential for harm to political stability, reputation, and public safety Marchal et al. (2024). Scholars have warned of an “infopocalypse” where constant exposure to fake media breeds cynicism or paranoia Schick (2020). Detecting deepfakes reliably is therefore critical to mitigate the spread of misinformation and disinformation [^1] We adopt the term misinformation throughout this paper to refer broadly to harmful or misleading content. Technically, misinformation denotes false information shared without intent to deceive, while disinformation refers to deliberately deceptive content. Our usage includes both, given the difficulty of inferring intent in large-scale media analysis., and to restore trust in digital media. The rapid advancement of AI-generated image technologies has reached a point where distinguishing between real and synthetic images has become increasingly challenging for humans. Studies have shown that humans underperform in identifying AI-generated images, highlighting the sophistication of these generative models (Diel et al., 2024).
@@
The political sphere is particularly vulnerable to the risks posed by deepfakes, which can be weaponized to manipulate public opinion and undermine democratic processes (Bengio et al., 2025c, b; Karen Hao, 2019). Synthetic media have already been exploited for scams, blackmail, and targeted reputation sabotage, while the fabrication of fake historical artifacts, manipulated medical images, and staged events introduces new avenues for the spread of misinformation and societal harm (Ferrara, 2024; Bengio et al., 2025c). By flooding social and traditional media with convincing falsehoods, deepfakes erode public trust in news and create confusion about what is real, particularly during sensitive periods like elections (IVADO and CEIMIA, 2025). Such disruptions threaten not only individual reputations and public safety, but also the legitimacy of democratic institutions and processes, with scholars warning that advances in generative AI could empower malicious actors to influence political outcomes and destabilize societies (Bengio et al., 2025a; Hameleers et al., 2024; Vaccari and Chadwick, 2020).
@@
Based on this political motivation, we built our dataset by starting with a large-scale collection of politically relevant social media posts, including tweets, images, and videos scraped during the 2025 Canadian federal election. While this initial focus helped identify image types likely to appear in high-impact or misleading contexts, the final dataset spans a broad range of content beyond political imagery, including everyday scenes, people, and objects. These posts serve as a foundation for analyzing real-world misinformation risks and curating image content with societal relevance. However, relying directly on social media data introduces challenges such as label contamination and unreliable ground truth, which we sought to avoid in contrast to prior work (Chen and Zou, 2023). To assess the deception risk of current generative models, we also conducted a user study measuring people’s ability to distinguish real from synthetic political images, revealing that outputs from some proprietary models may approach chance-level accuracy.
@@
Despite significant progress in deepfake detection research, current datasets suffer from major limitations that restrict their effectiveness in modern, real-world scenarios. As shown in Table˜1, most established benchmarks rely on outdated generation methods; GAN-based face-swapping tools such as DeepFaceLab (Liu et al., 2023) and Face2Face (Thies et al., 2016). These datasets, while valuable for early detection efforts, are increasingly unrepresentative of the latest synthetic media, particularly high-fidelity diffusion and transformer-based models. Moreover, they overwhelmingly focus on single-face portraits, providing little to no real-world grounded images and neglecting the broader spectrum of image-based misinformation that floods political and social media discourse: crowd scenes, protests, disaster images, manipulated signage, or synthetic screenshots.
To address these gaps, we introduce OpenFake [^2] https://huggingface.co/datasets/ComplexDataLab/OpenFake, a politically grounded dataset designed for general deepfake detection. OpenFake focuses on synthetic media that poses genuine misinformation risks. It pairs large-scale real image corpora with state-of-the-art synthetic images. A high-quality dataset like this helps benchmark existing detectors and supports the development of stronger ones. Generative and detection models evolve in a feedback loop: as new generative methods are released, new detectors must be built, which can then inform on directions to improve the generators even further. As generation methods advance rapidly, static datasets risk becoming outdated. To ensure our dataset remains relevant over time, we introduce OpenFake Arena, a crowdsourced adversarial platform where users are invited to generate synthetic images that fool a live deepfake detection model. Successful submissions are validated using a CLIP-based prompt similarity gate, scored against a continually updated classifier, and, if they pass, contributed back into the dataset, forming a self-improving benchmark. This enables continual learning and ensures the benchmark remains both current and increasingly challenging.
In summary, our contributions are:
@@
  • A large, politically relevant dataset of 3 million real images paired with extracted prompts, curated for misinformation risk and motivated by social media content.
@@
  • A diverse, high-quality synthetic image set spanning 963k samples, generated from both open-source and proprietary models.
  • A human perception study revealing that some modern proprietary generators can reduce human classification accuracy to nearly random chance.
  • Baseline analysis shows detectors trained on OpenFake outperform those from older datasets against high-quality deepfakes, showing its value for modern detection.
@@
  • A scalable crowdsourcing framework (OpenFake Arena) for adversarial image generation, enabling continual community-driven benchmarking.
@@ 2 Related work
Despite the proliferation of generative models, existing deepfake datasets remain limited in realism, diversity, and accessibility. Early benchmarks such as FaceForensics++ (Rössler et al., 2019), Celeb-DF (Li et al., 2020), and DFDC (Dolhansky et al., 2020) rely on GAN-based face-swapping techniques and focus almost exclusively on single-person portrait videos. Even newer datasets such as ForgeryNet (He et al., 2021), OpenForensics (Le et al., 2021), and FFIW (Zhou et al., 2021) continue to emphasize face-centric detection, with limited variation in image content or generation method (Cheng et al., 2024; Chen et al., 2024b; Yan et al., 2024). More recent image datasets have started to incorporate diffusion-based generators (e.g., Stable Diffusion, DALL·E 2, Midjourney), as seen in Fake2M (Lu et al., 2023), DiffusionForensics (Wang et al., 2023), and GenImage (Zhu et al., 2023). However, these datasets still fall short in several ways. First, most rely on open-source models like SDv1.5 or SDv2.1 (Rombach et al., 2022), which, while important, do not match the visual fidelity of cutting-edge proprietary models such as Imagen 3 (Baldridge et al., 2024) or GPT Image 1 (OpenAI, 2025). As a result, they fail to represent the modern threat landscape posed by the most deceptive fakes. Second, many datasets lack real-world grounding. Image prompts are frequently abstract, artistic, or class-based (e.g., GenImage uses classes from ImageNet-1k (Deng et al., 2009)), failing to capture the multimodal misinformation strategies actually deployed online. Third, these datasets are static and infrequently updated, meaning they quickly become outdated as generation tools evolve. Fourth, prompts used for image generation are often withheld, making it difficult for others to reproduce, regenerate, or expand these datasets with future models. In contrast, we release a large bank of extracted prompts along with the images, which enables researchers to extend the dataset using alternative text-to-image models while preserving topical consistency. Finally, accessibility remains a persistent issue. Many datasets require downloading large zip archives via Google Drive or web links, making them difficult to integrate into new pipelines. In contrast, OpenFake is fully hosted on the HuggingFace Hub in streaming-friendly Parquet format, enabling scalable access and evaluation, which should help the community develop new detection tools. Table˜1 highlights these differences in model coverage, dataset scope, prompt extensibility, and access modality.
@@
Dataset Year # Fakes # Reals Extendable Content Scope Realism Access # Methods Most recent model
FaceForensics++ (Rössler et al., 2019) 2019 5K 1K  Narrow Low  Hard 4 Face2Face (2016)
Celeb-DF (Li et al., 2020) 2020 5K+ 590  Narrow Low  Hard 1 DeepFaceLab (2020)
DFDC (Dolhansky et al., 2020) 2020 100K+ 20K+  Narrow Low  Hard 8 DeepFaceLab (2020)
ForgeryNet (He et al., 2021) 2021 1.5M 1.5M  Narrow Low  Hard 15 DeepFaceLab (2020)
FFIW (Zhou et al., 2021) 2021 10K 10K  Narrow Low  Hard 3 DeepFaceLab (2020)
OpenForensics (Le et al., 2021) 2021 100K 100K  Narrow Low  Hard 3 GAN (2020)
DeepFakeFace (Song et al., 2023) 2023 90K 30K  Narrow Medium  Hard 3 SD v1.5 (2022)
Fake2M (Lu et al., 2023) 2023 2M 0  Moderate Medium Easy 3 SD v1.5 (2022)
DiffusionForensics (Wang et al., 2023) 2023 570K 140K  Moderate Medium  Hard 8 iDDPM (2021)
DMDetection (Corvi et al., 2023) 2023 200K 200K  Moderate Medium  Hard 3 DALL·E 2 (2022)
GenImage (Zhu et al., 2023) 2023 1.3M 1.33M  Moderate Good  Hard 5 Midjourney 5 (2023)
TWIGMA (Chen and Zou, 2023) 2023 800K 0  Rich Medium  Unavailable
DiffusionDeepfake (Bhattacharyya et al., 2024) 2024 100K 94K  Narrow Good  Hard 2 Midjourney (2024)
DF40 (Yan et al., 2024) 2024 1M+ 1.5K  Narrow Medium  Hard 40 PixArt-$\alpha$ (2024)
DiffusionFace (Chen et al., 2024b) 2024 600K 30K  Narrow Good  Hard 11 SD v2.1 (2022)
DiFF (Cheng et al., 2024) 2024 500K 23K  Narrow Good  Hard 13 Midjourney 5 (2023)
Semi-Truths (Pal et al., 2024) 2024 1.34M 26K  Moderate Good Easy 8 Stable Diffusion XL (2023)
@@
Table 1: Overview of public deepfake datasets. “# Fakes” and “# Reals” count individual media items (images or videos; units omitted for brevity). Content Scope:  Narrow (face-focused);  Moderate (diverse but limited);  Rich (broad, internet-like variety). Access:  Unavailable;  Hard (public but cumbersome);  Easy (clean, ready-to-use hosting). Realism: qualitative fidelity of synthetic content — Low, Medium, Good, High. Extendable: availability of prompts/metadata enabling seamless dataset expansion. OpenFake uniquely combines rich scope, high realism, large real sample count, easy access, and extensibility.
@@
Early detection approaches relied on convolutional neural networks trained on known forgery artifacts, such as blending boundaries or low-level inconsistencies (Afchar et al., 2018; Rössler et al., 2019; V and Joy, 2023). While effective in-domain, these models struggle to generalize across generation techniques (Ojha et al., 2023). As diffusion and transformer-based models reduce such artifacts, recent work has focused on semantic-level signals and frequency-domain cues (Durall et al., 2020; Liu et al., 2021a; Frank et al., 2020; Qian et al., 2020). CLIP-based detection (Cozzolino et al., 2024; Khan and Dang-Nguyen, 2024) has emerged as a promising direction, leveraging large-scale vision-language pretraining to improve robustness. Other advances include domain-adaptive feature learning (Shao et al., 2025; Jia et al., 2024), zero-shot detectors (Lin et al., 2024), and hybrid approaches that blend local artifact patterns with global semantic reasoning (LI et al., 2024; Zhou et al., 2024; Ma et al., 2025). Despite progress, the rapid pace of generative model development continues to outstrip detection capabilities, motivating adaptive benchmarks like OpenFake Arena to assess robustness in a dynamic, adversarial setting.
3 Empirical Motivation: Real-World Misinformation and Human Limits
3.1 Social media content
Social media platforms, notably X (formerly Twitter), have become critical channels for political discourse, and consequently, for amplifying deepfake disinformation. To better understand the real-world impact of deepfakes, we conducted an analysis of images shared on X.

### v2

# OpenFake: An Open Dataset and Platform Toward Real-World Deepfake Detection

@@ Abstract
Deepfakes, synthetic media created using advanced AI techniques, pose a growing threat to information integrity, particularly in politically sensitive contexts. This challenge is amplified by the increasing realism of modern generative models, which our human perception study confirms are often indistinguishable from real images. Yet, existing deepfake detection benchmarks rely on outdated generators or narrowly scoped datasets (e.g., single-face imagery), limiting their utility for real-world detection. To address these gaps, we present OpenFake, a large politically grounded dataset specifically crafted for benchmarking against modern generative models with high realism, and designed to remain extensible through an innovative crowdsourced adversarial platform that continually integrates new hard examples. OpenFake comprises nearly four million total images: three million real images paired with descriptive captions and almost one million synthetic counterparts from state-of-the-art proprietary and open-source models. Detectors trained on OpenFake achieve near-perfect in-distribution performance, strong generalization to unseen generators, and high accuracy on a curated in-the-wild social media test set, significantly outperforming models trained on existing datasets. Overall, we demonstrate that with high-quality and continually updated benchmarks, automatic deepfake detection is both feasible and effective in real-world settings.
@@ 1 Introduction
Deepfakes, realistic synthetic media generated by AI, have emerged as a serious threat to the information ecosystem [9, 7]. By enabling anyone to fabricate audio-visual content of real people, deepfakes can spread false information at an unprecedented scale, eroding trust across various platforms, from social media and online content to traditional media outlets. High-profile cases (e.g., forged speeches or imagery of public figures) and the prevalence of non-consensual intimate imagery underscore the potential for harm to political stability, reputation, and public safety [52]. Scholars have warned of an “infopocalypse” where constant exposure to fake media breeds cynicism or paranoia [65]. Detecting deepfakes reliably is therefore critical to mitigate the spread of misinformation and disinformation [^1] We adopt the term misinformation throughout this paper to refer broadly to harmful or misleading content. Technically, misinformation denotes false information shared without intent to deceive, while disinformation refers to deliberately deceptive content. Our usage includes both, given the difficulty of inferring intent., and to restore trust in digital media. The rapid advancement of AI-generated image technologies has reached a point where distinguishing between real and synthetic images has become increasingly challenging for humans. Studies have shown that humans underperform in identifying AI-generated images, highlighting the sophistication of these generative models [19].
@@
The political sphere is particularly vulnerable to the risks posed by deepfakes, which can be weaponized to manipulate public opinion and undermine democratic processes [7, 6, 36]. Synthetic media have already been exploited for scams, blackmail, and targeted reputation sabotage, while the fabrication of fake historical artifacts, manipulated medical images, and staged events introduces new avenues for the spread of misinformation and societal harm [24, 7]. By flooding social and traditional media with convincing falsehoods, deepfakes erode public trust in news and create confusion about what is real, particularly during sensitive periods like elections [34]. Such disruptions threaten not only individual reputations and public safety, but also the legitimacy of democratic institutions and processes, with scholars warning that advances in generative AI could empower malicious actors to influence political outcomes and destabilize societies [5, 28, 71].
@@
Based on this political motivation, we built our dataset by starting with a large-scale collection of politically relevant social media posts, including tweets, images, and videos scraped during the 2025 Canadian federal election. While this initial focus helped identify image types likely to appear in high-impact or misleading contexts, the final dataset spans a broad range of content beyond political imagery, including everyday scenes, people, and objects. These posts serve as a foundation for analyzing real-world misinformation risks and curating image content with societal relevance. However, relying directly on social media data introduces challenges such as label contamination and unreliable ground truth, which we sought to avoid in contrast to prior work [12]. To assess the deception risk of current generative models, we also conducted a user study measuring people’s ability to distinguish real from synthetic political images, revealing that outputs from some proprietary models may approach chance-level accuracy.
@@
Despite significant progress in deepfake detection research [42, 67, 17], current datasets suffer from major limitations that restrict their effectiveness in modern, real-world scenarios [59, 12]. As shown in Table˜1, most established benchmarks rely on outdated generation methods; GAN-based face-swapping tools such as DeepFaceLab [46] and Face2Face [69]. These datasets, while valuable for early detection efforts, are increasingly unrepresentative of the latest synthetic media, particularly high-fidelity diffusion and transformer-based models. Moreover, they overwhelmingly focus on single-face portraits, providing little to no real-world grounded images and neglecting the broader spectrum of image-based misinformation that floods political and social media discourse: crowd scenes, protests, disaster images, manipulated signage, or synthetic screenshots.
To address these gaps, we introduce OpenFake [^2] https://huggingface.co/datasets/ComplexDataLab/OpenFake , a politically grounded dataset for general deepfake detection. OpenFake pairs large-scale real image corpora with state-of-the-art synthetic images and is designed to remain extensible through OpenFake Arena, a crowdsourced adversarial platform that continually contributes hard, validated examples via a CLIP-based prompt-consistency gate and scoring against a live detector. This yields a self-improving benchmark that tracks the evolution of modern generators.
By training a modest SwinV2-Small [48] detector on OpenFake we produce near perfect in-distribution results on the held-out test sets (unseen images from the seen generators and their variants), as well as strong performance on images from unseen generators (see Table˜3), with overall F1 score of 0.99 compared to the 0.88 from our strongest baseline (same model trained on GenImage). More interestingly, the model trained on OpenFake achieves a strong performance on a curated in-the-wild social-media test set, with a F1 score of 0.86 compared to 0.08 (GenImage), and 0.26 (Semi-Truths) (see Table˜4). This finding stands in contrast to the prevailing pessimism around automatic deepfake detection, which is deemed as both futile and intractable, largely due to the increasing realism of generative models [30, 23, 38, 40, 10, 2]. Indeed, some experts argue that distinguishing real from synthetic images will soon become impossible, which has shifted attention to watermarking [74, 55, 47, 64]. However, watermarking depends on developer cooperation, consistent deployment across proprietary systems, and robustness to post-processing, and therefore cannot replace open-world detection. OpenFake paves the way for automatic detection, demonstrating that highly realistic deepfake, which evade detection by human eyes, can be detected with high accuracy. In summary, our contributions are:
  • Providing OpenFake, an easily accessible dataset, which has
@@
    • Rich Scope: A large, politically relevant dataset of 3 million real images paired with extracted prompts, curated for misinformation risk and designed by studying real-world social media.
@@
    • High Realism: A diverse, high-quality synthetic image set spanning 963k images, generated from state-of-the-art open-source and proprietary models.
@@
    • Extendable: A scalable crowdsourcing framework (OpenFake Arena) for adversarial image generation, enabling continual community-driven benchmarking.
  • An experimental study which shows the weak performance of detectors trained on the currently available datasets on detecting realistic deepfakes, with F1 ranging between 0.0 to 0.88. Our model trained on OpenFake significantly outperform these baselines with a F1 of 0.99.
  • A human perception study showing that images from modern proprietary generators can be imperceptible to humans, with accuracy in some cases dropping to near random chance (e.g., with Imagen 3). Our model trained on OpenFake, however, achieves a near perfect performance.
  • A real-world feasibility study which report a strong performance of OpenFake on detecting real and fake images actually circulated on social media, based on a carefully curated in-the-wild test set. OpenFake achieves a F1 score of 0.86, significantly higher that when relying on two strong contenders: 0.08 (GenImage), and 0.26 (Semi-Truths).
@@ 2 Related work
Despite the proliferation of generative models, existing deepfake datasets remain limited in realism, diversity, and accessibility. Early benchmarks such as FaceForensics++ [63], Celeb-DF [43], and DFDC [20] rely on GAN-based face-swapping techniques and focus almost exclusively on single-person portrait videos. Even newer datasets such as ForgeryNet [29], OpenForensics [41], and FFIW [78] continue to emphasize face-centric detection, with limited variation in image content or generation method [15, 13, 77]. More recent image datasets have started to incorporate diffusion-based generators (e.g., Stable Diffusion, DALL·E 2, Midjourney), as seen in Fake2M [50], DiffusionForensics [73], and GenImage [80]. However, these datasets still fall short in several ways. First, most rely on open-source models like SDv1.5 or SDv2.1 [62], which, while important, do not match the visual fidelity of cutting-edge proprietary models such as Imagen 3 [4] or GPT Image 1 [58]. As a result, they fail to represent the modern threat landscape posed by the most deceptive fakes. Second, many datasets lack real-world grounding. Image prompts are frequently abstract, artistic, or class-based (e.g., GenImage uses classes from ImageNet-1k [18]), failing to capture the multimodal misinformation strategies actually deployed online. Third, these datasets are static and infrequently updated, meaning they quickly become outdated as generation tools evolve. Fourth, prompts used for image generation are often withheld, making it difficult for others to reproduce, regenerate, or expand these datasets with future models. In contrast, we release a large bank of extracted prompts along with the images, which enables researchers to extend the dataset. Finally, accessibility remains a persistent issue. Many datasets require downloading large zip archives via Google Drive or web links, making them difficult to integrate into new pipelines. In contrast, OpenFake is fully hosted on the HuggingFace Hub in streaming-friendly Parquet format, enabling scalable access and evaluation, which should help the community develop new detection tools. Table˜1 highlights these differences in model coverage, dataset scope, prompt extensibility, and access modality.
@@
Dataset Year Fakes Reals Extends Content Scope Realism Access Methods Most recent model
FaceForensics++ [63] 2019 5K 1K Narrow Low  Hard 4 Face2Face (2016)
Celeb-DF [43] 2020 5K+ 590 Narrow Low  Hard 1 DeepFaceLab (2020)
DFDC [20] 2020 100K+ 20K+ Narrow Low  Hard 8 DeepFaceLab (2020)
ForgeryNet [29] 2021 1.5M 1.5M Narrow Low  Hard 15 DeepFaceLab (2020)
FFIW [78] 2021 10K 10K Narrow Low  Hard 3 DeepFaceLab (2020)
OpenForensics [41] 2021 100K 100K Narrow Low  Hard 3 GAN (2020)
DeepFakeFace [68] 2023 90K 30K Narrow Medium  Hard 3 SD v1.5 (2022)
Fake2M [50] 2023 2M 0 Moderate Medium Easy 3 SD v1.5 (2022)
DiffusionForensics [73] 2023 570K 140K Moderate Medium  Hard 8 iDDPM (2021)
DMDetection [16] 2023 200K 200K Moderate Medium  Hard 3 DALL·E 2 (2022)
GenImage [80] 2023 1.3M 1.33M Moderate Good  Hard 5 Midjourney 5 (2023)
TWIGMA [12] 2023 800K 0 Rich Medium  Unavailable
DiffusionDeepfake [8] 2024 100K 94K Narrow Good  Hard 2 Midjourney (2024)
DF40 [77] 2024 1M+ 1.5K Narrow Medium  Hard 40 PixArt-$\alpha$ (2024)
DiffusionFace [13] 2024 600K 30K Narrow Good  Hard 11 SD v2.1 (2022)
DiFF [15] 2024 500K 23K Narrow Good  Hard 13 Midjourney 5 (2023)
Semi-Truths [59] 2024 1.34M 26K Moderate Good Easy 8 Stable Diffusion XL (2023)
@@
Table 1: Compared to current public deepfake datasets, OpenFake uniquely combines rich scope, high realism, large real sample count, easy access, and extensibility. “Fakes” and “Reals” count individual media items (images or videos; units omitted for brevity). Content Scope: Narrow (face-focused); Moderate (diverse but limited); Rich (broad, internet-like variety). Access: Unavailable; Hard (public but cumbersome); Easy (clean, ready-to-use hosting). Realism: qualitative fidelity of synthetic content — Low, Medium, Good, High. Extendable: availability of prompts/metadata enabling seamless dataset expansion.
@@
Early detection approaches relied on convolutional neural networks trained on known forgery artifacts, such as blending boundaries or low-level inconsistencies in the images [1, 63, 70]. While effective in-domain, these models struggle to generalize across generation techniques [56]. As diffusion and transformer-based models reduce such artifacts, recent work has focused on semantic-level signals and frequency-domain cues [21, 45, 25, 60]. CLIP-based detection [17, 37] has emerged as a promising direction, leveraging large-scale vision-language pretraining to improve robustness. Other advances include domain-adaptive feature learning [67, 35], zero-shot detectors [44], and hybrid approaches that blend local artifact patterns with global semantic reasoning [42, 79, 51]. Despite progress, the rapid pace of generative model development continues to outstrip detection capabilities, motivating adaptive benchmarks like OpenFake Arena to assess robustness in a dynamic, adversarial setting.
3 Case Study: Real-World Misinformation and Human Limits